# One-Shot Method for Computing Generalized Winding Numbers

### v1

@@ 5.1 Implementation Details
To implement this generic algorithm on a particular geometry representation, we need three main procedures: finding intersections between boundary curves projected onto the unit sphere, ray-surface intersection, and computing areas on the sphere.
Meshes
For meshes, we find ray-mesh intersections using a k-d tree; areas are computed using the spherical polygon area formula as simply $A=2\pi-\sum_{i}(\alpha_{i}-\pi),$ where $\alpha_{i}$ are the interior angles. Finally, we identify regions by computing a surface arrangement on the sphere by first finding intersections between polylines on a sphere, which can be done using a Surface Sweep Algorithm. For a maximum constant number of intersection points, its time complexity is $\mathcal{O}(B\log B)$ Berg et al. (2008), where $B$ is the number of mesh boundary segments. All in all, the asymptotic time complexity of our algorithm is $\mathcal{O}(B\log B+\log F)$ for a single query point. For such scenarios as voxelization, however, the ray intersection is done only one for many query points, so the time complexity for meshes is $\mathcal{O}(B\log B+\frac{1}{\max(W,H)}\log F)$ per point, which for meshes with $B\ll F$ is faster than the state-of-the-art exact winding number computation with the empirical complexity of $\mathcal{O}(F^{0.55})$ Jacobson et al. (2013). Note that while Barill et al. (2018) computes a much faster approximation, our algorithm computes exact winding number.
Parametric Surfaces
For surfaces made out of parametric patches, such as Coons patches, Bézier patches or triangles, or NURBS patches, we can compute the winding number within machine precision without surface discretization as follows. Starting with the problem of ray-parametric patch intersection, let’s take a Bézier triangle as an example, which in essence is a low-degree polynomial $f(u,v)$ with a simplex $u,v>0,u+v<1$ as the parametric domain.
@@
As we show in the Supplementary materials, this, as well as the problems of intersecting two parametric curves, such as Bézier, and finding their self-intersections, can be relaxed into a Sum-of-Squares (SOS) formulations Marschner et al. (2021), which then yield convex problems solvable by standard semidefinite programming (SDP) solvers. This relaxation is tight, yielding exact recovery in almost all cases (Sec. 7, Marschner et al. (2021)). All these techniques, as shown in Marschner et al. (2021) are easily extendable to arbitrary NURBS patches.
Finally, for a region bounded by a parametric curve $\gamma(s):[0,L]\rightarrow S_{p}^{2}$ projected onto the unit sphere, using Gauss-Bonnet theorem, we can compute the area inside the curve as $A=2\pi-\int_{0}^{L}k_{g}(s)$, where $k_{g}$ is the geodesic curvature of the boundary. Note that for most parametric boundaries, including Bézier curves, this integral has to be evaluated numerically. Unlike surface discretization, however, this boundary discretization may lead only to minor differences in the final winding number close to the boundary (Sec. 7).
Following this logic, in practice we can discretize the boundary of a parametric surface and compute region decomposition using polyline intersection algorithm, just like in the mesh case. All the results in the paper are computed via in that manner for optimal performance.
Once the arrangement on the sphere is computed, whether for parametric curves or polylines, we find a single point in the largest region Wein et al. (2024).
Speed up
@@
SOS optimization in general can be slow (see Sec. 7). If speed is more important than precision, in practice we can find ray-surface intersections using a nonlinear equation solver. In this case, we can express the ray-surface intersection test as an equation $f(\xi)+(q-p)t=0$, which we solve numerically for $t\geq 0,\xi\in\Omega$, finding all roots. We first can compute an axis-aligned bounding box a parametric patch via SOS Marschner et al. (2021), then intersect it with each ray, giving $t_{min}$ and $t_{max}$; box’s diagonal is $diag$. If the parametric surface has no self-intersections, we then compute all roots by dividing the interval $[t_{\rm{min}},t_{\rm{max}}]$ into a number of subintervals (in our implementation, we use $\max(2,30\frac{t_{\rm{max}}-t_{\rm{min}}}{diag})$ and initializing the solver with the middle of each subinterval. We then consider the roots duplicate, if the distance between them is less than the accuracy of the solver. If surface can have self-intersections, we recommend using SOS optimization for accuracy.
@@ 6 Curve Networks
In this section, we demonstrate an application of our algorithm to determining winding number of a curve network, where each patch satisfies a known linear PDE. In the following, we focus on the Laplace’s equation for the simplicity of exposition; our method can be generalized to any PDE where Boundary Element Method is applicable, such as Poisson equation, biharmonic equation, or others.
@@ 6.1 Preprocessing
Given an input curve network composed of polylines, we use the method of Zhuang et al. (2013) to detect loops, consisting of input curve segments. We then identify boundary as edges adjacent to only one loop. If the surface is an orientable 2-manifold, we then consistently orient the loops via breadth-first search such that each non-boundary curve segment is traversed once in each direction. This procedure detects if the surface is not a manifold, in which case we compute a winding number as a sum of winding numbers of each manifold patch.
@@ 6.2 Ray-Surface Intersection
$\begin{split}&\Delta x=0\\ &\left.x\right|_{\partial\Omega}=\tilde{x},\end{split}$ (5)
where $\tilde{x}$ is the given curve loop $x$ coordinate. For simplicity, for the examples we set $\Omega=[0,1]^{2}$. Solutions to this problem are known to be minimal surfaces.
Here we follow the Boundary Element Method (BEM) with collocation LaForce (2006). We express the surface as a smooth map $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, where each coordinate of $f(\xi)=(x,y,z)$ is the solution to the corresponding boundary value problem (Eq. 5). In BEM, a representation formula states that the solution of the linear differential equation at a point $\xi\in\Omega$ is expressible as a boundary integral over $\Gamma=\partial\Omega$, which converts our surface into parametric form:
@@
where $G(\xi,\eta)$ is the fundamental solution of the PDE, and $n$ is the outwards normal of $\Gamma$ (see Supplementary for details).
We discretize the integrals in Eq. 6 using constant boundary elements, and numerically integrate for each polyline segment using trapezoid integration. To find Neumann boundary conditions, i.e., the normal derivative $\left.\frac{\partial f(\xi)}{\partial n}\right|_{\Gamma}$, we use the Dirichlet boundary conditions, which are the curve loop’s coordinates, and solve a dense linear system LaForce (2006) via Cholesky decomposition.
Once both boundary conditions are known, we apply our method using the representation formula in Eq. 6 as a parametric form. We further optimize the process by observing that a minimal surface is contained inside a bounding box of the boundary $\tilde{x}$ due to the maximum principle. We can therefore easily compute the minimum and maximum possible $t$ values $t_{\rm{min}},t_{\rm{max}}$ by intersecting the ray with the axis-aligned bounding box, improving performance.
@@ 7 Validation and Results
We implemented our method in C++ with MATLAB subroutines for SOS optimization. We used Eigen Guennebaud et al. (2010), YALMIP Löfberg (2004) as the SOS interface and MOSEK ApS (2024) to solve. We use sphericalpolygon library Li (2020) for areas. We use GNU Scientific Library’s Galassi et al. (1996) implementation of Powell’s Hybrid method as a nonlinear solver when not using SOS.
@@
Throughout the paper, we demonstrate the results of our one-shot winding algorithm on various parametric surfaces (a set of Coons patches in Fig. 1a and 14b, extrusion surfaces in Fig. 2d,14a), meshes (Fig. 1c, 13), and curve networks with minimal surfaces (Fig. 1b, 11). We demonstrate surfaces with boundary, because without boundary our method is a trivial inside-outside test. The boundaries on the curve networks are selected manually. In Fig. 12, we show a typical application of winding numbers, a voxelization of the volume enclosed inside the curve network, computed without discretizing the surfaces. We consider a voxel inside if its center has a winding number $\geq 0.5$. We demonstrate boolean operations in Fig. 15. To compute the ‘set union’, we add the winding numbers; to compute ‘set intersection’, we multiply them pointwise.
SOS accuracy
We validate the accuracy of the SOS solvers, showing all the intersections with parametric surfaces can be found within machine precision. Note that Bézier curves and Bézier triangles were chosen as an example; Marschner et al. (2021); Zhang et al. (2023) demonstrate that all these techniques are easily generalizable to arbitrary NURBS patches.
In each test, we generate the random control points, as well as ray origin and direction, following a uniform $[-1,1]$ distribution. We generated 400 test pairs (half with intersections, half without) for each test. For the ray intersections, we got exact recovery (within machine precision) or correctly identified no intersections for $99.75\%$ of the rays. Exact recovery is identified as a corresponding eigenvalue being larger than a threshold ($10^{-3}$ in our experiments) Marschner et al. (2021). The only ray that did not have exact recovery had an error in $t$ of 0.0064; such rare ray can be simply ignored using the eigenvalue threshold. The projected Bézier curve intersection test has a similar exact recovery rate of 99.75%, same for self-intersections.
Parametric with discretized boundary
@@
We validate the accuracy of our one-shot winding number computation for a discretized boundary and compare it with the accuracy of the hierarchical winding number algorithm (HWN) Jacobson et al. (2013) and Barill et al. (2018) that use a mesh (Fig. 9). As as simple test case, we took a randomly generated Bézier triangle, generated $10^{4}$ query points evenly in an arbitrary slice, and plotted maximum and mean $\ell_{2}$ (RMSE) errors in a logarithmic scale of our method and theirs as functions of the number of boundary samples and mesh vertices, respectively. For the mesh-based methods, we meshed the parametric domain with a regular triangle mesh. The horizontal axis indicates the number of boundary edges for our method and their mesh (bottom), and the corresponding number of faces (top).
@@
Even with a small number of boundary edges, our method is quite precise. Both mesh-based methods, until a certain refinement level, have large maximum errors close to the surface: Each time a query point is on one side of the parametric surface and on the other of the mesh, the winding number will have a wrong sign, yielding a significant error. In contrast, the error of our method, even for roughly discretized boundaries, is significantly smaller, noticeable only around the surface boundary, and stems mainly from approximation of the spherical areas. Using SOS completely eliminates all others sources of error.
@@
As refinement grows, FWN Barill et al. (2018), being an approximation, does not improve. Our method becomes an order of magnitude more precise than the exact mesh-based Jacobson et al. (2013); our worst (maximum) error is roughly the same as their average. Note that to get to a similar maximum error accuracy as our $1900$ boundary edges, mesh-based methods would need roughly $16.3\cdot 10^{6}$ faces, taking 0.9Gb in memory. Our memory footprint is negligible in comparison (kilobytes).
Performance
We evaluate performance of our method on a Intel® Core™ i9-7900X @ 3.30GHz. SOS, while being precise, is slower than our C++ code: finding intersections of projected Bézier curves takes $592\pm 275ms$ via SOS versus $1.2$ms via the polyline intersection code; ray-Bézier triangle intersection takes $581\pm 25$ms via SOS versus $8.85$ms via the nonlinear solver with our parameters. Most of SOS time is spent in YALMIP, not in the solver. We therefore measure all the total performances using the C++ code, parallelized via OpenMP.
@@
The bottleneck of our algorithm is marked as ’Boundary Processing’ (BP); almost all time is spent on computing finding all intersections between spherical segments. Note that despite the optimal algorithmic complexity of $O(B\log B)$, in our implementation we use a naïve $O(B^{2})$ algorithm which may explain the performance.
For meshes, the accuracy of our method is exactly the same as the ground truth of Jacobson et al. (2013). Performance-wise, our method may bring advantages for meshes where, as discussed in Sec. 5.1, the boundary has few elements, while the mesh has many faces. In Table 8 we present typical performance statistics for such meshes. Note that the approximation algorithm Barill et al. (2018) timing statistics are a few times faster: $0.59,0.55,0.53\mu$s. Our goal, however, is precise winding number computation: as Fig. 9 shows, their approximation is often imprecise.
@@
Figure 7: Timing statistics for curve networks with minimal surfaces (top half) and parametric surfaces with Coons patches (Car, Ship), and extrusion surfaces (gear, roll). The columns #$\#L,\#B,\#Q$ refer to loops, boundary edges, and query points respectively. Total time in ms per query point is in the last column; time spent boundary processing is %BP.
Input # F # B # Q % B. P. HGWN ($\mu$s) Ours ($\mu$s)
Piggybank 11k 6 600k 96.66 4.41 1.46
@@
Figure 8: Bottom: statistics for meshes, time in $\mu$s ($10^{-6}s$) per query point. HGWN refers to the hierarchical method of Jacobson et al. (2013).
Curve networks with minimal surfaces
@@
We compare the accuracy of the nonlinear solver approach for a BEM curve network, with the method of Jacobson et al. (2013) on a FEM discretization of the surface using linear elements. As the ground truth surface, we take $z(x,y)=e^{x-1}\sin(y)-e^{x}\cos(y)$; we mesh it and compute its winding numbers. We then reconstruct the surface by FEM using a regular grid on the parametric domain. Then we uniformly sample $342$k points, keeping only those 2368 points with the absolute value of the ground truth winding number $\in(0.4995,0.5005)$; on the rest the error is 0. We can see (Fig. 10) that our solution is as accurate as theirs for just a fraction of the stored data (e.g., 500 vertices vs. 5000 triangles). The discrete increments of the error are because each query point can be classified as either on the correct side of the surface, yielding zero error, and on the wrong side, yielding a winding number error of 1.
Limitations
When a surface has multiple boundaries, such as polygon soups, our assumption of constant number of boundary intersections when projected onto the sphere is no longer true, and the method becomes less efficient.
When not using SOS, the accuracy of our computations relies on the nonlinear equation solver that we use. Errors in root finding, especially missed or spurious roots, can change the value of $\chi$, causing errors. While our analysis (Sec. 7) shows that this rare, for precision applications one should use SOS.

### v2

@@ 5.1 Implementation Details
To implement this generic algorithm, we need three main procedures: finding intersections between boundary curves projected onto the unit sphere, ray-surface intersection, and computing areas on the sphere.
Meshes. For meshes, we find ray-mesh intersections using a k-d tree; areas are computed using the spherical polygon area formula as simply $A=2\pi-\sum_{i}(\alpha_{i}-\pi),$ where $\alpha_{i}$ are the interior angles. Finally, we identify regions by computing a surface arrangement on the sphere by first finding intersections between polylines on a sphere, which can be done using a Surface Sweep Algorithm. For a maximum constant number of intersection points, its time complexity is $\mathcal{O}(B\log B)$ [BCKO08], where $B$ is the number of mesh boundary segments. All in all, the asymptotic time complexity of our algorithm is $\mathcal{O}(B\log B+\log F)$ for a single query point. For such scenarios as voxelization, however, the ray intersection is done only once for many query points, so the time complexity for meshes is $\mathcal{O}(B\log B+\frac{1}{\max(W,H)}\log F)$ per point, which for meshes with $B\ll F$ is faster than the state-of-the-art exact winding number computation with the empirical complexity of $\mathcal{O}(F^{0.55})$ [JKS13]. Note that while [BDS18] computes a much faster approximation, our algorithm computes exact winding number.
Figure 8: The results of our one-shot algorithm on meshes. Camel head mesh [Sor05], Vase by virtox.
Parametric Surfaces. For surfaces made out of parametric patches, such as Coons patches, Bézier patches or triangles, or NURBS patches, we can compute the winding number without surface discretization as follows. Starting with the problem of ray-parametric patch intersection, let us take a Bézier triangle as an example, which in essence is a low-degree polynomial $f(u,v)$ with a simplex $u,v>0,u+v<1$ as the parametric domain.
@@
Figure 9: Additional results on parametric surfaces: a spiral with high winding numbers (a) and a non-manifold boat with Coons patches (b)
@@
In practice, we use a nonlinear solver to find intersections. We express the ray-surface intersection test as an equation $f(\xi)+(q-p)t=0$, which we solve numerically for $t\geq 0,\xi\in\Omega$, finding all roots. We first can compute an axis-aligned bounding box of a parametric patch, then intersect it with each ray, giving $t_{min}$ and $t_{max}$; box’s diagonal is $D$. If the parametric surface has no self-intersections, we then compute all roots by dividing the interval $t_{\textrm{min}},t_{\textrm{max}}]$ into a number of subintervals (in our implementation, we use $\max(2,30\frac{t_{\textrm{max}}-t_{\textrm{min}}}{D})$ and initializing the solver with the middle of each subinterval. We then consider the roots duplicate, if the distance between them is less than the accuracy of the solver. Note that for polynomial patches one can use specialized techniques such as ‘pencil of a matrix’ [XBC19]. We do not assume that surfaces are polynomial, so in our 3D implementation we use a generic nonlinear solver instead.
We discretize the boundary of a parametric surface and compute region decomposition using polyline intersection algorithm, similar to the mesh case. Unlike surface discretization, however, this boundary discretization may lead only to minor differences in the final winding number close to the boundary (Sec. 7).
To find a point within a spherical region, we first compute the Euclidean centroid of the region, which, due to the convexity of the sphere, lies inside the sphere. We project it onto the sphere (inset, white circle) and connect it to an arbitrary point (one of the black crosses) along the region boundary via a geodesic, i.e., great circle. We then choose the middle of some great circle arc that is inside the polygon (inset, white star), testing whether an arc is inside using the orientation of the boundary curve tangents at the arc’s endpoints. Note that by construction the circle will necessarily either intersect the region boundary or be tangent to it. If it is tangent, we choose another arbitrary boundary point and repeat the procedure.
Multiple Boundaries. In the case that the surface has multiple disconnected boundaries, we project all the boundaries in a single spherical arrangement.
Exact Computations. For parametric geometry, our pipeline enables exact computation of the winding number. As we show in Appendix C, the intersection problem (Eq. 4), as well as the problems of intersecting two parametric curves, such as Bézier, and finding their self-intersections, can be relaxed into Sum-of-Squares (SOS) formulations [MZPS21], which then yield convex problems solvable by standard semidefinite programming (SDP) solvers. This relaxation is tight, yielding exact recovery in almost all cases (Sec. 7, [MZPS21]). All these techniques, as shown in [MZPS21] are easily extendable to arbitrary NURBS patches. In this case, for exactness, instead of sampling a point within the largest region on the sphere, we sample a random point on the sphere and perform a point location within the spherical region decomposition, which can be done exactly. SOS optimization, however, is slow (see Sec. 7). Finally, for a region bounded by a parametric curve $\gamma(s):[0,L]\rightarrow S_{p}^{2}$ projected onto the unit sphere, using Gauss-Bonnet theorem, we can compute the area inside the curve as $A=2\pi-\int_{0}^{L}k_{g}(s)$, where $k_{g}$ is the geodesic curvature of the boundary. Note that for most parametric boundaries, including Bézier curves, this integral has to be evaluated numerically.
@@ 6 Curve Networks
In this section, we demonstrate an application of our algorithm determining winding number of a curve network (Fig. 11, 18), where each patch satisfies a known linear PDE. In the following, we focus on the Laplace’s equation for the simplicity of exposition; our method can be generalized to any PDE where Boundary Element Method is applicable, such as Poisson equation, biharmonic equation, or others.
Figure 10: Curve networks can, depending on the context, represent different surfaces, such as (a) parametric patches [BWSS12], (b) minimal surfaces [WC21], [Web19] (c) other PDEs [NISA07], or (d) more complex surfaces [YAB22].
@@ 6.1 Preprocessing
Given an input curve network composed of polylines, we use the method of [ZZCJ13] to detect loops, consisting of input curve segments. We then identify boundaries as segments adjacent to only one loop. If the surface is an orientable 2-manifold, we then consistently orient the loops via breadth-first search such that each non-boundary curve segment is traversed once in each direction. This procedure detects if the surface is not a manifold, in which case we compute a winding number as a sum of winding numbers of each manifold patch.
@@ 6.2 Ray-Surface Intersection
[…]
where $\tilde{x}$ is the given curve loop $x$ coordinate. For simplicity, for the examples we set $\Omega=[0,1]^{2}$. Solutions to this problem are related to minimal surfaces [MP12].
Here we follow the Boundary Element Method (BEM) with collocation [LaF06]. We express the surface as a smooth map $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, where each coordinate of $f(\xi)=(x,y,z)$ is the solution to the corresponding boundary value problem (Eq. 5). In BEM, a representation formula states that the solution of the linear differential equation at a point $\xi\in\Omega$ is expressible as a boundary integral over $\Gamma=\partial\Omega$, which converts our surface into parametric form:
@@
where $G(\xi,\eta)$ is the fundamental solution of the PDE, and $n$ is the outwards normal of $\Gamma$ (See Appendix B).
We discretize the integrals in Eq. 6 using constant boundary elements, and numerically integrate for each polyline segment using trapezoid integration. To find Neumann boundary conditions, i.e., the normal derivative $\left.\frac{\partial f(\xi)}{\partial n}\right|_{\Gamma}$, we use the Dirichlet boundary conditions, which are the curve loop’s coordinates, and solve a dense linear system [LaF06] via Cholesky decomposition.
Once both boundary conditions are known, we apply our method using the representation formula in Eq. 6 as a parametric form. We further optimize the process by observing that a minimal surface is contained inside a bounding box of the boundary $\tilde{x}$ due to the maximum principle. We can therefore easily compute the minimum and maximum possible $t$ values $t_{\textrm{min}},t_{\textrm{max}}$ by intersecting the ray with the axis-aligned bounding box, improving performance.
@@ 7 Validation and Results
We implemented our method in C++ with Eigen [GJ10], sphericalpolygon library [Li20] for areas, and GNU Scientific Library’s [GDT96] implementation of Powell’s Hybrid method as a nonlinear solver for intersection with parametric geometry.
@@
Throughout the paper, we demonstrate the results of our one-shot winding algorithm on various 2D parametric curves, parametric surfaces (a set of Coons patches in Fig. One-Shot Method for Computing Generalized Winding Numbers a and 9b, extrusion surfaces in Fig. 1d,9a), meshes (Fig. One-Shot Method for Computing Generalized Winding Numbersc, 8), and curve networks with minimal surfaces (Fig. One-Shot Method for Computing Generalized Winding Numbersb, 18). We demonstrate surfaces with boundary, because without boundary our method is a trivial inside-outside test. The boundaries on the curve networks are selected manually. In Fig. 11, we show a typical application of winding numbers, a voxelization of the volume enclosed inside the curve network, computed without discretizing the surfaces. We consider a voxel inside if its center has a winding number $\geq 0.5$. We demonstrate boolean operations in Fig. 12. To compute the ‘set union’, we add the winding numbers; to compute ‘set intersection’, we multiply them pointwise.
2D Parametric Geometry. We first evaluate the performance and precision of our method on 2D parametric geometry. As a test, we randomly generate cubic Bézier curves (N=1000) with control points in the unit square. For each curve, we compute a winding number for each point of a regular grid ($250\times 250$) in the curve’s bounding box (Fig. 13).
Figure 13: For 2D parametric curves, on a typical regular grid of query points, our method (a) is more precise and faster than the state-of-the-art method [SGW24] (b). The performance of both methods is minimally affected by the stopping criterion threshold, since most of the computation time is spent calculating arccosines. We find that our method is $4\times$ faster on a regular grid of $250\times 250$ query points within the bounding box of 1000 random cubic Bézier curves. Both methods fall back to the same algorithm when the query point is outside the bounding box of the curve.
We compare our performance with the 2D state-of-the-art method [SGW24]. For any query point outside the bounding box, we use the same closure property as [SGW24], so our performance on those points is identical. Inside the bounding box, on a regular grid of query points, however, our method performs significantly faster (up to $4\times$), as we only need to perform subdivisions for one ray per row due to our one-shot strategy. Furthermore, in this setup, our algorithm is more precise: We obtain 0 ‘misclassified’ points (winding number error above $10^{-10}$ compared to the ground truth) with a threshold of $10^{-7}$. Our method takes 5074ms,while [SGW24] takes 20131ms and produces 47 misclassified points.
[…]
@@
Parametric Surfaces. We validate the accuracy of our one-shot winding number computation for a discretized boundary and compare it with the accuracy of the hierarchical winding number algorithm (HWN) [JKS13] and [BDS18] that use a mesh (Fig. 15). As as simple test case, we took a randomly generated Bézier triangle, generated $10^{4}$ query points evenly in an arbitrary slice, and plotted maximum and mean $\ell_{2}$ (RMSE) errors in a logarithmic scale of our method and theirs as functions of the number of boundary samples and mesh vertices, respectively. For the mesh-based methods, we meshed the parametric domain with a regular triangle mesh. The horizontal axis indicates the number of boundary edges for our method and their mesh (bottom), and the corresponding number of faces (top).
@@
Even with a small number of boundary edges, our method is quite precise. Both mesh-based methods, until a certain refinement level, have large maximum errors close to the surface: Each time a query point is on one side of the parametric surface and on the other of the mesh, the winding number will have a wrong sign, yielding a significant error. In contrast, the error of our method, even for roughly discretized boundaries, is significantly smaller, noticeable only around the surface boundary, and stems mainly from approximation of the spherical areas.
@@
Under refinement, FWN [BDS18], being an approximation, does not improve. Our method becomes an order of magnitude more precise than the exact mesh-based [JKS13]; our worst (maximum) error is roughly the same as their average. Note that to get to a similar maximum error accuracy as our $1900$ boundary edges, mesh-based methods would need roughly $16.3\cdot 10^{6}$ faces, taking 0.9Gb in memory. Our memory footprint is negligible in comparison (kilobytes).
Figure 15: Logarithmic error plot with respect to ground truth for our method (blue), [JKS13] (HGWN, red) and [BDS18] (FWN, green). Even when the boundary is discretized, our method computes winding numbers a few orders of magnitude more precisely than [JKS13] or [BDS18] on a mesh of comparable complexity. Horizontal axis shows the number of faces (top) and boundary vertices (bottom) of a mesh and used in our method.
3D Performance. We evaluate performance of our method on a Intel® Core™ i7-9700X @ 4.900GHz. We measure all the total performances using the C++ code, parallelized via OpenMP.
@@
The bottleneck of our algorithm is marked as ‘Boundary Processing’ (BP); almost all time is spent on computing finding all intersections between spherical segments. Note that despite the optimal algorithmic complexity of $O(B\log B)$, in our implementation we use a naïve $O(B^{2})$ algorithm which may explain the performance.
[…]
@@
[…]
Piggy 11k 6 600k 96.66 4.41 1.46
@@
[…]
@@
Curve networks with minimal surfaces. We compare the accuracy of the nonlinear solver approach for a BEM curve network, with the method of [JKS13] on a FEM discretization of the surface using linear elements. As the ground truth surface, we take $z(x,y)=e^{x-1}\sin(y)-e^{x}\cos(y)$; we mesh it and compute its winding numbers. We then reconstruct the surface by FEM using a regular grid on the parametric domain. Then we uniformly sample 342342k points, keeping only those 2368 points with the absolute value of the ground truth winding number $\in(0.4995,0.5005)$; on the rest the error is 0. We can see (Fig. 14) that our solution is as accurate as theirs for just a fraction of the stored data (e.g., 500 vertices vs. 5000 triangles). The discrete increments of the error are because each query point can be classified as either on the correct side of the surface, yielding zero error, and on the wrong side, yielding a winding number error of 1.
Adaptive Quadrature. For parametric curves and surfaces, the generalized winding number can be computed via numerically evaluating the signed solid angle surface integral using adaptive quadrature. In 2D, for a regular grid of query points, we are significantly faster than the state-of-the-art method [SGW24], which was already an improvement over adaptive quadrature. In 3D, our approach is generally slower than adaptive quadrature methods, with the bottleneck being the computation of spherical arrangement.
For a class of parametric surfaces, however, with a simpler boundary, for instance, planar, our method is faster in a scenario with a regular grid. As a projection of a planar curve on a sphere can have no self-intersections, for such parametric surfaces we only need to perform a trivial area computation and a ray-surface intersection. We show two examples of surfaces of revolution with planar boundaries in Fig. 16 where we are both faster and more accurate than adaptive quadrature. On a regular grid of $800\times 800$, it is 83s for our method vs. 103s for adaptive quadrature (both yielded similar errors of 3.1e-5, Fig. 16a). Since the parametric surface in Fig. 16b has a sharp feature, we switched both methods to higher precision arithmetic, decreasing performance. On the challenging example in Fig. 16b, our method computed the winding numbers in 484s with an MSE of 2.6e-4. Adaptive quadrature took 12637s to compute result with a high MSE of 0.0809. This showcases the critical difference with the adaptive quadrature: since our method does not need to discretize the surface, high-frequency surface details are much less of a problem for us. Our approach also requires significantly less function evaluations compared to adaptive quadrature.
Figure 16: Our approach on two surfaces of revolutions. Computing the winding numbers for the surface in (a) took our method 83s while adaptive quadrature took 103s for a similar error of 3.125e53.125e-5 and 3.140e53.140e-5 respectively. In (b), as the surface has a sharp feature, we used higher precision arithmetic, significantly decreasing performance. For this experiment, our method took 484s while having a MSE of 2.6875e42.6875e-4. This was an almost intractable input for adaptive quadrature, taking it 12637s to only achieve a MSE of 0.0809.
Singular and unbounded surfaces. Our method is compatible with both singular and unbounded surfaces, as long as the boundary has an asymptote at infinity and each ray has a finite number of intersections. We demonstrate an example in Fig. 17. For unbounded surfaces, we project the asymptote of the surface at $\lim_{r\rightarrow\infty}$ from the center of the parametric domain, otherwise our method requires no changes. Note that these surfaces are hard to process with conventional methods, as discretizing them may result in significant accuracy loss.
Figure 17: Our method works without modification on any oriented surfaces that support ray-intersection queries, including infinite surfaces with singularities.
Precise computation for parametric surfaces. We also implemented a proof of concept SOS optimization for precise winding number computation for parametric shapes, in MATLAB. We use YALMIP [Löf04] as the SOS interface and MOSEK [ApS24] to solve.
We validate the accuracy of the SOS solvers, showing all the intersections with parametric surfaces can be found within machine precision. Note that Bézier curves and Bézier triangles were chosen as an example; [MZPS21, ZMST23] demonstrate that all these techniques are easily generalizable to arbitrary NURBS patches.
In each test, we generate the random control points, as well as ray origin and direction, following a uniform [1,1][-1,1] distribution. We generated 400 test pairs (half with intersections, half without) for each test. For the ray intersections, we got exact recovery (within machine precision) or correctly identified no intersections for $99.75\%$ of the rays. Exact recovery is identified as a corresponding eigenvalue being larger than a threshold ($10^{-3}$ in our experiments) [MZPS21]. The only ray that did not have exact recovery had an error in $t$ of 0.0064; such rare ray can be simply ignored using the eigenvalue threshold. The projected Bézier curve intersection test has a similar exact recovery rate of 99.75%, same for self-intersections.
SOS, while being precise, is slow: finding intersections of projected Bézier curves takes $592\pm 275ms$ via SOS versus 1.21.2ms via the polyline intersection code; ray-Bézier triangle intersection takes $581\pm 25$ms via SOS versus 8.858.85ms via the nonlinear solver with our parameters. Most of SOS time is spent in YALMIP, not in the solver.
Limitations. As the complexity of our method is driven by the complexity of the boundary, our method excels when the boundary is simple, and is slow when the boundary is complex or has multiple connected components. Perhaps an efficient system would choose a suitable method to compute the winding number based on the boundary complexity vs. the complexity of the curve/surface itself, falling back to previous methods if necessary. Additionally, errors in root finding, especially missed or spurious roots, can change the value of $\chi$, causing errors. While our analysis (Sec. 7) shows this is extremely rare, this can be fixed by increasing the search resolution or, when exactness is necessary, via SOS.